Known vectors:

⃗a = (1, 0, 1) , ⃗b = (0, 1, -1) , ⃗c = (0, 0, 1)

Find the angle between:

1. a and b

2. a and c

3. b and c


1
Expert's answer
2021-11-26T11:58:28-0500
"\vec a \cdot \vec b=1(0)+0(1)+1(-1)=-1"

"\vec a \cdot \vec c=1(0)+0(0)+1(1)=1"

"\vec b \cdot \vec c=0(0)+1(0)-1(1)=-1"

"|\vec a|=\sqrt{1^2+0^2+1^2}=\sqrt{2}"

"|\vec b|=\sqrt{0^2+1^2+(-1)^2}=\sqrt{2}"

"|\vec c|=\sqrt{0^2+0^2+1^2}=1"

1.

"\cos \angle(\vec a, \vec b)=\dfrac{\vec a \cdot \vec b}{|\vec a||\vec b|}=\dfrac{-1}{\sqrt{2}(\sqrt{2})}=-\dfrac{1}{2}"

"\angle(\vec a, \vec b)=120\degree"

2.

"\cos \angle(\vec a, \vec c)=\dfrac{\vec a \cdot \vec c}{|\vec a||\vec c|}=\dfrac{1}{\sqrt{2}(1)}=\dfrac{\sqrt{2}}{2}"

"\angle(\vec a, \vec c)=45\degree"

3.


"\cos \angle(\vec b, \vec c)=\dfrac{\vec b \cdot \vec c}{|\vec b||\vec c|}=\dfrac{-1}{\sqrt{2}(1)}=-\dfrac{\sqrt{2}}{2}"

"\angle(\vec b, \vec c)=135\degree"